40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1683     Entry time: Sun Jul 24 20:03:38 2016
Author: Antonio 
Type: DailyProgress 
Category: South Cavity 
Subject: Mode matching and cavities locking 

Today I have aligned the South cavity with the beam prepared and described in the previous ID entry.

The South cavity is aligned and mode matched with a visibility (Vmax-Vmin)/Vmax of ~75%.


Things done today:

  • Aligned the beam in order to get the TEM00 mode resonating into the cavity;
    The resulting visibility is of 0.75;
  • Prepared and connected all the power supplies to their components for the south FSS lock;
  • South Cavity locked; The lock happened this morning and it is still locked; (1mW Pin).
  • Realigned (vis = ~ 0.7) and reset all the polarization in the North path, because for some unknown reason
    the North cavity was not anymore in the same place of two days ago, not even close; the waveplates too.
  • Worked on the locking of the North Cavity as it was not showing robust; The North cavity stays locked for few hours now.
  • Started to preapare the Optics for the PLL:
    - Photodiode for the beat note, mirrors and beam splitters alignment;
    - Mirrors and beam splitters alignment;
    - Alignment of the transmitted beams from the cavities into the beat note PD together with Andrews
      The two beams are visually overlapped now;
    - We checked the polarizations at which the two cavities are locked currently:
    North is locked on P;
    South mostly on S (not totally);
    - I and Andrew decided to use 2 lambda/2 at the output in order to be able to change the linear polarization whenever
    we need. The optics at the output are mainly made for s-light;
  • Photodiodes and optics for the ISS are aligned too;
  • I have connected the power supply for the temperature (output=10V);
  • Started some tyding up the table;


We can start to look at the beat note!

ELOG V3.1.3-